OpenStudy (anonymous):

If u and v are the vectors below, find the vector w whose tail is at the point halfway from the tip of v to the tip of u and whose head is at the point halfway from the tip of u to the tip of u-v. u = (-2,-4,-5) v = (-3, 5, -1) answer: (2, -7, -3/2)

OpenStudy (anonymous):

To find halfway between u and v, calculate u-v and divide by 2. To find halway from u to u-v, calculate u-(u-v) and divide by 2. Subtract the two points you get and there's your answer!

OpenStudy (anonymous):

ok ill try it, thanks!

OpenStudy (anonymous):

when you say divide by 2, do you mean multiply each number by 1/2?

OpenStudy (anonymous):

The answer is (2,-7, -3/2) how do i get it?

OpenStudy (anonymous):

u = (-2,-4,-5) v = (-3, 5, -1)I don't know if I'm too late... (u-v) = (1,-9,-4) P=1/2(u-v) = (1/2,-9/2,-2) u-(u-v) = (-3,5,-1) Q=1/2(u-(u-v))=(-3/2,5/2,-1/2) P-Q=(2,-7,-3/2) Make sense?